OpenStudy (anonymous):

NEED HELP!!! PLEASE?!!?!!!! This question says to use Elimination to solve the problem and enter it in form like this (4, 5). I am very horrible at math and have had a few tutors try to help me with it and I still cannot understand. Need help. Problem I have is this: 2x + 6y = -8 5x - 3y = 88

OpenStudy (anonymous):

Well, let's start with the basics. An equation is essentially an equality (it has the "=" sign) between one or more variables (x,y,z,q, any letter of the alphabet is used to represent it). 2x=4 is an equation which describes the equality between 2x and 4. By "solving" we generally mean that we are interested in what x is equal to, 2x=4 is solved by dividing the whole equation by 2 to end up with x=2 which is our answer here. This is allowed because it is one of the fundamental properties of equations. Let me go a bit in depth with this. Let's say an apple is worth 2$ or apple=2$. Would it be true to say that 2 apples=4$? Or that 3 apples = 6$? Or that 10 apples=20$? Sure thing! Here we can deduce that: * if (something)=(another thing) then (a number) * (something) = (the same number) * (another thing) Same thing goes for division: if an apple=2$ then half an apple is 1$, 1/4 of an apple is 50 cents and so forth. * if (something)=(another thing) then (something) / (a number) = (another thing) / (the same number) What about addition and subtraction ? If apple=2$ would it be true that apple+3$=5$? Of course it would! We can further deduce that: * if (something)=(another thing) then (a number) + (something) = (the same number) + (another thing) Similarly for subtraction: * if (something)=(another thing) then (something) - (a number) = (another thing) - (the same number)

OpenStudy (anonymous):

I still don't get it... I'm sorry. Is there any way you could do an example problem for me?